# LLM-Powered Social Digital Twins: A Framework for Simulating Population Behavioral Response to Policy Interventions

### v1

@@ Abstract
Predicting how populations respond to policy interventions is a fundamental challenge in computational social science and public policy. Traditional approaches rely on aggregate statistical models that capture historical correlations but lack mechanistic interpretability and struggle with novel policy scenarios. We present a general framework for constructing Social Digital Twins virtual population replicas where Large Language Models (LLMs) serve as cognitive engines for individual agents. Each agent, characterized by demographic and psychographic attributes, receives policy signals and outputs multi-dimensional behavioral probability vectors. A calibration layer maps aggregated agent responses to observable population-level metrics, enabling validation against real-world data and deployment for counterfactual policy analysis.
We instantiate this framework in the domain of pandemic response, using COVID-19 as a case study with rich observational data. On a held-out test period, our calibrated digital twin achieves a 20.7% improvement in macro-averaged prediction error over gradient boosting baselines across six behavioral categories. Counterfactual experiments demonstrate monotonic and bounded responses to policy variations, establishing behavioral plausibility. The framework is domain-agnostic: the same architecture applies to transportation policy, economic interventions, environmental regulations, or any setting where policy affects population behavior. We discuss implications for policy simulation, limitations of the approach, and directions for extending LLM-based digital twins beyond pandemic response.
@@ 1 Introduction
Traditional approaches to policy prediction fall into two categories. Aggregate statistical models (econometric regressions, time series forecasting) capture historical correlations but lack mechanistic interpretability: they can say what happened but not why [1]. When policies are novel or contexts shift, these models extrapolate poorly. Agent-based models (ABMs) offer mechanistic richness by simulating individual decision-makers [2, 3], but require extensive manual specification of decision rules a knowledge bottleneck that limits applicability to domains where behavior is well-understood.
Recent advances in Large Language Models (LLMs) suggest a new paradigm. LLMs trained on vast corpora of human-generated text have learned implicit models of human reasoning, preferences, and decision-making [4, 5]. Researchers have demonstrated that LLMs can reproduce survey responses when conditioned on demographic attributes [6], engage in realistic social interactions [7], and serve as proxies for human subjects in economic experiments [8]. These capabilities suggest using LLMs as cognitive engines for agent-based simulation replacing hand-coded decision rules with neural models that have learned human-like reasoning.
@@ 1.1 The Digital Twin Paradigm
    Validation Protocol: Comparison against real-world observational data with strict temporal separation
This architecture is domain-agnostic. The same framework can simulate:
@@
The key insight is that LLMs have learned general models of human decision-making from training data spanning all these domains. By conditioning on appropriate context, we can instantiate domain-specific behavioral simulators without domain-specific training.
@@ 1.2 Case Study: COVID-19 Pandemic Response
    Rich observational data: Google Mobility Reports [10] and survey trackers [11] provide high-frequency behavioral measurements
@@
    Multi-dimensional behavior: Six mobility categories (work, retail, grocery, transit, parks, residential) capture diverse behavioral adaptations
@@ 1.3 Contributions
    A General Framework: We formalize Social Digital Twins as LLM-powered agent simulations with calibrated output mappings, applicable across policy domains
@@
    Multi-Dimensional Behavioral Modeling: Unlike prior work extracting scalar compliance probabilities, our agents output multi-dimensional behavioral vectors enabling richer validation
@@
    Calibration Methodology: We present a multi-objective optimization approach for learning agent-to-observation mappings from historical data
@@
    Rigorous Evaluation Protocol: We demonstrate strict train/test separation, per-category metrics, counterfactual sanity checks, and ablation studies establishing component necessity
The remainder of this paper is organized as follows: Section 2 reviews related work across agent-based modeling, LLM simulation, and digital twins. Section 3 presents the general framework. Section 4 details our pandemic response instantiation and experimental results. Section 5 discusses implications, limitations, and extensions. Section 6 concludes.
@@ 2 Related Work
2.1 Agent-Based Modeling for Policy
Agent-based models (ABMs) simulate populations as collections of autonomous agents following decision rules [2]. In economics, ABMs model market dynamics [24]; in epidemiology, disease transmission [25]; in urban planning, transportation choices [26]. Classical ABMs require explicit rule specification, typically based on utility maximization [27] or behavioral heuristics [28]. This knowledge bottleneck limits applicability: modelers must anticipate relevant decision factors and encode them explicitly.
@@ 2.2 LLMs as Behavioral Simulators
These works establish that LLMs contain useful behavioral priors. However, most lack calibration against observational data they evaluate face validity or alignment with controlled experiments rather than real-world prediction accuracy.
@@ 2.3 Digital Twins
The digital twin concept originated in manufacturing as virtual replicas of physical systems [13]. Extensions to urban systems [14], healthcare [15], and social systems [16] have been proposed. Policy-oriented digital twins aim to simulate citizen responses to interventions, but most implementations use simplified behavioral models (e.g., gravity models for mobility, discrete choice for mode selection) rather than LLM-driven reasoning.
@@ 2.4 Retrieval and Grounding for LLMs
Parallel work on improving LLM factuality through retrieval-augmented generation [17], knowledge editing [18], and attention biasing [19] is conceptually related. These methods ground LLM outputs in external knowledge; our calibration layer similarly grounds agent outputs in observational data. The key difference is that we ground behavioral predictions rather than factual claims.
@@ 2.5 Positioning This Work
  • Multi-agent simulation (from ABM tradition)
@@
Unlike prior LLM simulation work, we calibrate against real-world observational data with strict temporal holdout. Unlike traditional ABMs, we replace hand-coded rules with LLM-generated behavioral outputs. Unlike aggregate forecasting, we maintain individual-level heterogeneity and can perform counterfactual analysis.
@@ 3.1 Framework Overview
Figure 1: Social Digital Twin Framework. The architecture is domain-agnostic: policy signals, agent attributes, and observable outputs are instantiated per-domain, while the LLM engine and calibration methodology are general.
@@ 3.3 LLM Cognitive Engine
Each dimension represents the probability of engaging in a specific behavior (e.g., going to work, using public transit, wearing a mask). The multi-dimensional output is crucial: it enables validation against diverse observational signals and captures behavioral trade-offs (e.g., less transit use $\rightarrow$ more car use).
@@ 3.4 Aggregation
Individual agent outputs are aggregated to population-level predictions. For simple cases, we use the mean:
@@ 3.5 Calibration Layer
In our implementation, we use per-dimension linear mappings with clipping:
@@
We use Optuna [20] with Tree-structured Parzen Estimator for multi-objective optimization, balancing errors across behavioral dimensions.
@@ 3.6 Validation Protocol
    Per-Dimension Metrics: Report accuracy for each behavioral category, not just aggregate
@@
    Baseline Comparison: Compare against non-agentic alternatives (statistical models, persistence)
@@ 3.7 Counterfactual Analysis
Unlike black-box forecasting, we can ask: “What would behavior have been under different policies?” Such counterfactuals are not causally identified from observational data alone [21], but they provide behavioral plausibility checks. If the twin shows reasonable responses to policy variations (e.g., stricter policy $\rightarrow$ more compliance), we gain confidence in its internal reasoning.
@@ 4 Case Study: Pandemic Response
We instantiate the framework for COVID-19 pandemic response in the UAE.
4.1 Domain-Specific Configuration
@@ 4.3.2 Analysis
  • Workplaces (+89.3%): LLM captures policy-work relationship (“lockdown $\rightarrow$ work from home”)
  • Retail (+37.8%), Parks (+32.6%): Discretionary activities are policy-sensitive; LLM reasoning outperforms statistical extrapolation
@@ 5.1 Implications for Policy Simulation
Our results suggest LLM-powered digital twins are viable for policy simulation, with specific strengths and limitations:
@@ 6 Conclusion
We present a general framework for Social Digital Twins LLM-powered simulations of population behavioral response to policy interventions. The framework comprises synthetic agent populations with demographic attributes, LLM cognitive engines generating multi-dimensional behavioral outputs, calibration layers grounding predictions in observational data, and rigorous validation protocols.

### v2

# LLM Powered Social Digital Twins: A Framework for Simulating Population Behavioral Response to Policy Interventions

@@ Abstract
Predicting how populations respond to policy interventions is a fundamental challenge in computational social science and public policy. Traditional approaches rely on aggregate statistical models that capture historical correlations but lack mechanistic interpretability and struggle with novel policy scenarios. We present a general framework for constructing Social Digital Twins virtual population replicas where Large Language Models (LLMs) serve as cognitive engines for individual agents. Each agent, characterized by demographic and psychographic attributes, receives policy signals and outputs multi dimensional behavioral probability vectors. A calibration layer maps aggregated agent responses to observable population level metrics, enabling validation against real world data and deployment for counterfactual policy analysis.
We instantiate this framework in the domain of pandemic response, using COVID-19 as a case study with rich observational data. On a held out test period, our calibrated digital twin achieves a 20.7% improvement in macro averaged prediction error over gradient boosting baselines across six behavioral categories. Counterfactual experiments demonstrate monotonic and bounded responses to policy variations, establishing behavioral plausibility. The framework is domain agnostic: the same architecture applies to transportation policy, economic interventions, environmental regulations, or any setting where policy affects population behavior. We discuss implications for policy simulation, limitations of the approach, and directions for extending LLM based digital twins beyond pandemic response.
@@ 1 Introduction
Traditional approaches to policy prediction fall into two categories. Aggregate statistical models (econometric regressions, time series forecasting) capture historical correlations but lack mechanistic interpretability: they can say what happened but not why [1]. When policies are novel or contexts shift, these models extrapolate poorly. Agent based models (ABMs) offer mechanistic richness by simulating individual decision makers [2, 3], but require extensive manual specification of decision rules a knowledge bottleneck that limits applicability to domains where behavior is well understood.
Recent advances in Large Language Models (LLMs) suggest a new paradigm. LLMs trained on vast corpora of human generated text have learned implicit models of human reasoning, preferences, and decision making [4, 5]. Researchers have demonstrated that LLMs can reproduce survey responses when conditioned on demographic attributes [6], engage in realistic social interactions [7], and serve as proxies for human subjects in economic experiments [8]. These capabilities suggest using LLMs as cognitive engines for agent based simulation replacing hand coded decision rules with neural models that have learned human like reasoning.
@@ 1.1 The Digital Twin Paradigm
    Validation Protocol: Comparison against real world observational data with strict temporal separation
This architecture is domain agnostic. The same framework can simulate:
@@
The key insight is that LLMs have learned general models of human decision making from training data spanning all these domains. By conditioning on appropriate context, we can instantiate domain specific behavioral simulators without domain specific training.
@@ 1.2 Case Study: COVID-19 Pandemic Response
    Rich observational data: Google Mobility Reports [10] and survey trackers [11] provide high frequency behavioral measurements
@@
    Multi dimensional behavior: Six mobility categories (work, retail, grocery, transit, parks, residential) capture diverse behavioral adaptations
@@ 1.3 Contributions
    A General Framework: We formalize Social Digital Twins as LLM powered agent simulations with calibrated output mappings, applicable across policy domains
@@
    multi dimensional Behavioral Modeling: Unlike prior work extracting scalar compliance probabilities, our agents output multi dimensional behavioral vectors enabling richer validation
@@
    Calibration Methodology: We present a multi objective optimization approach for learning agent to observation mappings from historical data
@@
    Rigorous Evaluation Protocol: We demonstrate strict train/test separation, per category metrics, counterfactual sanity checks, and ablation studies establishing component necessity
The remainder of this paper is organized as follows: Section 2 reviews related work across agent based modeling, LLM simulation, and digital twins. Section 3 presents the general framework. Section 4 details our pandemic response instantiation and experimental results. Section 5 discusses implications, limitations, and extensions. Section 6 concludes.
@@ 2 Related Work
2.1 Agent Based Modeling for Policy
Agent based models (ABMs) simulate populations as collections of autonomous agents following decision rules [2]. In economics, ABMs model market dynamics [24]; in epidemiology, disease transmission [25]; in urban planning, transportation choices [26]. Classical ABMs require explicit rule specification, typically based on utility maximization [27] or behavioral heuristics [28]. This knowledge bottleneck limits applicability: modelers must anticipate relevant decision factors and encode them explicitly.
@@ 2.2 LLMs as Behavioral Simulators
These works establish that LLMs contain useful behavioral priors. However, most lack calibration against observational data they evaluate face validity or alignment with controlled experiments rather than real world prediction accuracy.
@@ 2.3 Digital Twins
The digital twin concept originated in manufacturing as virtual replicas of physical systems [13]. Extensions to urban systems [14], healthcare [15], and social systems [16] have been proposed. Policy oriented digital twins aim to simulate citizen responses to interventions, but most implementations use simplified behavioral models (e.g., gravity models for mobility, discrete choice for mode selection) rather than LLM driven reasoning.
@@ 2.4 Retrieval and Grounding for LLMs
Parallel work on improving LLM factuality through retrieval augmented generation [17], knowledge editing [18], and attention biasing [19] is conceptually related. These methods ground LLM outputs in external knowledge; our calibration layer similarly grounds agent outputs in observational data. The key difference is that we ground behavioral predictions rather than factual claims.
@@ 2.5 Positioning This Work
  • Multi agent simulation (from ABM tradition)
@@
Unlike prior LLM simulation work, we calibrate against real world observational data with strict temporal holdout. Unlike traditional ABMs, we replace hand coded rules with LLM generated behavioral outputs. Unlike aggregate forecasting, we maintain individual level heterogeneity and can perform counterfactual analysis.
@@ 3.1 Framework Overview
Figure 1: Social Digital Twin Framework. The architecture is domain agnostic: policy signals, agent attributes, and observable outputs are instantiated per domain, while the LLM engine and calibration methodology are general.
@@ 3.3 LLM Cognitive Engine
Each dimension represents the probability of engaging in a specific behavior (e.g., going to work, using public transit, wearing a mask). The multi dimensional output is crucial: it enables validation against diverse observational signals and captures behavioral trade offs (e.g., less transit use $\rightarrow$ more car use).
@@ 3.4 Aggregation
Individual agent outputs are aggregated to population level predictions. For simple cases, we use the mean:
@@ 3.5 Calibration Layer
In our implementation, we use per dimension linear mappings with clipping:
@@
We use Optuna [20] with Tree structured Parzen Estimator for multi objective optimization, balancing errors across behavioral dimensions.
@@ 3.6 Validation Protocol
    Per Dimension Metrics: Report accuracy for each behavioral category, not just aggregate
@@
    Baseline Comparison: Compare against non agentic alternatives (statistical models, persistence)
@@ 3.7 Counterfactual Analysis
Unlike black box forecasting, we can ask: “What would behavior have been under different policies?” Such counterfactuals are not causally identified from observational data alone [21], but they provide behavioral plausibility checks. If the twin shows reasonable responses to policy variations (e.g., stricter policy $\rightarrow$ more compliance), we gain confidence in its internal reasoning.
@@ 4 Case Study: Pandemic Response
We instantiate the framework for COVID 19 pandemic response in the UAE.
4.1 Domain Specific Configuration
@@ 4.3.2 Analysis
  • Workplaces (+89.3%): LLM captures policy work relationship (“lockdown $\rightarrow$ work from home”)
  • Retail (+37.8%), Parks (+32.6%): Discretionary activities are policy sensitive; LLM reasoning outperforms statistical extrapolation
@@ 5.1 Implications for Policy Simulation
Our results suggest LLM powered digital twins are viable for policy simulation, with specific strengths and limitations:
@@ 6 Conclusion
We present a general framework for Social Digital Twins LLM powered simulations of population behavioral response to policy interventions. The framework comprises synthetic agent populations with demographic attributes, LLM cognitive engines generating multi dimensional behavioral outputs, calibration layers grounding predictions in observational data, and rigorous validation protocols.